Ticket Number: HEA-NOC/20070924-12 Ticket Status: UPDATE
Ticket Type: unscheduled Resolver: HEA-NOC
Ticket Opened: 20070924 17:37 UTC+1 Problem Start: 20070924 16:27 UTC+1
Ticket Update: 20070925 15:07 UTC+1 Problem End:
Site/line: An Cheim
an-cheim-sw1 and dcu-sw1 began emitting storm control filtering messages at
approx. 16:27. Investigation shows the percentage of detected broadcast
traffic has gone over the threshold of 2%.
Storm control appears to have put the following interfaces into a
- dcu-sw1, Gi2/0/4 (Tony Ryan Student Link)
- an-cheim-sw1 Gi1/0/2 (Bonded uplink 2 to cwt-sw1 headend)
20070924 17:15 RG Open ticket to track work.
20070925 15:05 RG Storm control syslog messages stopped at 17:35:20
yesterday. Will keep ticket open for investigation.
This ticket can be monitored at http://www.hea.net/tickets/20070924-12
HEAnet Limited. Registered in Ireland, No. 275301.